OpenStudy (lovelyharmonics):

Find the rectangular coordinates of the point with the polar coordinates (8, 3 divided by 2 pi).

OpenStudy (anonymous):

i really hope it is \((8,\frac{2\pi}{3})\) and not what you wrote

OpenStudy (anonymous):

not "3 divided by 2 pi" but rather "2 pi divided by 3" am i right?

OpenStudy (lovelyharmonics):

noooo im sorry but its \[(8,\frac{ 3 }{ 2 }\pi)\]

OpenStudy (anonymous):

ok that works too, 3 pi divided by 2

OpenStudy (anonymous):

it is always true that \((r,\theta)\) has rectangular form \(x=r\cos(\theta), y = r\sin(\theta)\) in your case it will be \[x=8\cos(\frac{3\pi}{2}),y=8\sin(\frac{3\pi}{2})\] do you know how to evaluate those?

OpenStudy (lovelyharmonics):

no .-.

OpenStudy (anonymous):

do you know \(\cos(\frac{3\pi}{2})\)? if the answer is "no" i will show you how to find it

OpenStudy (lovelyharmonics):

can you show me :D

OpenStudy (anonymous):

find \(\frac{3\pi}{2}\) on the unit circle on the last page of the attached cheat sheet then look at the coordinates the first coordinate is cosine, the second coordinate is sine

OpenStudy (anonymous):

let me know when you see it and what you get

OpenStudy (lovelyharmonics):

so 270degrees so its 0,-1?

OpenStudy (anonymous):

yes

OpenStudy (lovelyharmonics):

so could the answer be 0,-8

OpenStudy (anonymous):

that means \(x=8\times 0=0\) and \(y=8\times -1=-8\) so your point is \((0,-8)\)

OpenStudy (anonymous):

yes

OpenStudy (lovelyharmonics):

i beat you to it c: can i ask so more questions and tag you in them?

OpenStudy (anonymous):

yes